What is the intersection of the quadratic equations -x^2+1 and x^2-2x-3
Answer by stanbon(75887)

What is the intersection of the quadratic equations -x^2+1 and x^2-2x-3
When they intersect they must be equal:
----
Equation:
-x^2+1 = x^2-2x-3
2x^2 - 2x -4 = 0
x^2 -x - 2 = 0
Factor:
(x-2)(x+1) = 0
x = 2 or x = -1
------
If x = 2, y = -(2^2)+1 = -3
If x = -1, y = -(-1)^2 + 1 = 2
-----
Solution: (2,-3) and (-1,2)
===============================
Cheers,
Stan H.